# A Recipe for CAC: Mosaic-based Generalized Loss for Improved Class-Agnostic Counting

### v1

@@ Abstract
Class agnostic counting (CAC) is a vision task that can be used to count the total occurrence number of any given reference objects in the query image. The task is usually formulated as a density map estimation problem through similarity computation among a few image samples of the reference object and the query image. In this paper, we point out a severe issue of the existing CAC framework: Given a multi-class setting, models don’t consider reference images and instead blindly match all dominant objects in the query image. Moreover, the current evaluation metrics and dataset cannot be used to faithfully assess the model’s generalization performance and robustness. To this end, we discover that the combination of mosaic augmentation with generalized loss is essential for addressing the aforementioned issue of CAC models to count objects of majority (i.e. dominant objects) regardless of the references. Furthermore, we introduce a new evaluation protocol and metrics for resolving the problem behind the existing CAC evaluation scheme and better benchmarking CAC models in a more fair manner. Besides, extensive evaluation results demonstrate that our proposed recipe can consistently improve the performance of different CAC models. The code will be released upon acceptance.
@@ 1 Introduction
Through in-depth analysis, we pinpoint the root causes of RO problem: the lack of variety of classes in each query image during the training stage incapacitates the model’s need to differentiate among object classes. Additionally, using pixel-wise losses (e.g., MSE) fails to accurately localize target objects and still contributes to the model’s inability to differentiate objects effectively. To address these challenges, we propose the Mosaic-based Generalized-Loss (MG) recipe for CAC, which consistently allows different CAC models to effectively capture variations in reference appearances and precisely localize target objects as illustrated in Figure 1(b) and demonstrated by our extensive experimental results. Finally, we also provide a strong CAC baseline by exploiting more advanced cross-attention modules from [5] as the backbone with our proposed MG loss recipe for training, denoted as MGCAC. The proposed MGCAC achieves the state-of-the-art results for CAC upon the original FSC-147 [17] and our proposed mosaic evaluation datasets.
@@ 2 Related Works
Due to a large number of related works in the literature, we briefly review the relevant ones on class-specific and class-agnostic object counting as follows.
@@ 3.3 Benchmarking
Furthermore, the current CAC evaluation protocol evaluates model counting performance by aggregating counting errors across all images. Given groundtruth count $c_{l}$ and predicted count $\tilde{c_{l}}$ for the $l$th image in the dataset with $L$ images, $\text{MAE}=\frac{1}{L}\sum_{i=1}^{L}|\tilde{c_{l}}-c_{l}|$ and $\text{RMSE}=\sqrt{\frac{1}{L}\sum_{i=1}^{L}(\tilde{c_{l}}-c_{l})^{2}}$ are usually used. The challenge arises from the long-tailed distribution of per-image object counts, introducing bias towards large-count query images. Figure 4 illustrates that only two images featuring large counts in FSC-147 test dataset, with 1123.jpg standing out as the outlier with the highest object count. As shown in Figure 6, the model prediction errors on 1123.jpg have notable squared error discrepancies. By excluding the first two highest object-count query images from RMSE calculations on the FSC-147 test dataset (Table 1(b)), we show significant impacts on the evaluation results. Recently SOTA models (i.e., LOCA and CounTR) get worse RMSE than BMNet+ when removing large-count queries. Those phenomena reinforce the need for an evaluation strategy that accounts for the distributional characteristics of object counts, especially in datasets with varying count sizes.
@@
Figure 6: Comparison of CAC models evaluated on 1123.jpg from FSC-147 dataset. Both BMNet+ and CounTR exhibit a considerable difference in squared error compared to LOCA.
@@ 4.1 Mosaic Evaluation Dataset
Currently, there is a lack of real-world multi-class counting datasets for CAC evaluation. We leverage the FSC-147 dataset as a foundation to create a multi-class setting. As shown in Figure 5, for a synthesized mosaic collage, we randomly select 4 query images of distinct classes from the validation and test sets of FSC-147. These images are then randomly cropped to the size of 384 × 384 and put together as size 768 × 768. Each class in the collage will individually serve as the reference class. Given the associated reference images, the model has to accurately predict the number of target objects for the collage (i.e. predicted count should equate to the ground truth count of the query image containing the target class). In total, we generate 1,000 queries (i.e., 4,000 data pairs), facilitating a comprehensive evaluation of the object recognition capabilities within the context of mosaic collages.
@@ 5.2 Generalized Loss
Currently, most CAC approaches based on density map estimation employ the pixel-wise $L_{2}$ loss (or MSE loss) to train the models where the ground truth density map is generated by converting the point annotations of each training image through convolving with a Gaussian kernel. However, the counting performance is sensitive to the hyperparameter selection of the Gaussian kernel since it greatly affects the generated ground truth density map. Moreover, $L_{2}$ loss focuses on pixel-level alignment while overlooking object-level localization. Recently, Ma et al. [14] propose a Generalized Loss (GL) that uses unbalanced optimal transport (OT) to measure the transport cost between the predicted density map and the ground truth point annotations directly. In addition, it also has shown that the $L_{2}$ loss is a special case of GL and usually results in suboptimal solutions. Thus, we mainly exploit GL to train the proposed models shown as follows:
@@ 6.1 Implementation Details
Datasets. We utilize three datasets to evaluate our model performance: (1) FSC-147 dataset [17], a standard class-agnostic object counting dataset composed of 147 categories and 6,135 images across different scales, in which each image is annotated with three reference objects. Given that the dataset contains decent annotations, we use the FSC-147 dataset to train and evaluate our models. (2) To ensure the generalizations of our models on unseen datasets, we evaluate FSC-147-trained models on CARPK [7], a dataset that contains 1,448 photos of cars in parking lots from a bird-view perspective. (3) Moreover, we evaluate FSC-147-trained models on our proposed mosaic dataset to validate their generalizability in multi-class scenarios.
@@
Model Comparison. We utilize MixFormer [5] architecture design with a multi-scale density head to better localize features of different levels as our baseline, denoted as MGCAC as mentioned in Section 1. (Please refer to Appendix for more detailed descriptions.) It is worth noting that our training recipe can be applied to other CAC model architectures as well. To demonstrate the effectiveness and general applicability of our proposed training recipe, we compare BMNet+, CounTR, and LOCA with their original training approach against our proposed recipe on the mosaic dataset.
@@ 6.2 Evaluation Results with Other SOTAs
Mosaic Evaluation. Seen from Table 3, models under the CAC typical training process, such as BMNet+ and LOCA, exhibit unsatisfactory performance under mosaic evaluation. The introduction of mosaic augmentation, as demonstrated by CounTR, yields a better performance. After employing our training recipe, BMNet+, CounTR, and LOCA exhibit substantial error decreases in all metrics. Specifically, BMNet+ yields an improvement of 47.5% in RMSE and over 75% across the remaining three metrics. These improvements demonstrate the importance and efficacy of applying the combination of mosaic augmentation and generalized loss on CAC models in discriminating reference classes and accurately matching targets. Notably, MGCAC achieves SOTA on the mosaic dataset with 37% improvement on all metrics against all models, showing our superiority in counting target objects.
@@
Cross-Dataset Generalization. A strong CAC model should excel in counting objects on scenarios that differ from the training dataset in terms of perspectives, scales, illumination, etc. To evaluate the ability to generalize on other datasets, we train the proposed method on the FSC-147 dataset and then evaluate on the CARPK testing set without the use of fine-tuning on the CARPK training set. Specifically, since the CARPK dataset mainly consists of car images, we remove the car category from the FSC-147 dataset during the training stage. In Table 4, our models outperform the non-fine-tuned baselines, which demonstrates MGCAC’s generalizability in diverse counting scenarios.
Table 3: Evaluation results on the mosaic dataset on the FSC-147 dataset of the CAC models trained with the proposed training recipe (Red) against those using their original training recipe where we use their publicly available pre-trained models (Green).
@@
Table 4: Cross-dataset evaluation results compared with other SOTA models where ‘fine-tuned’ means the model is further fine-tuned using the CARPK training data.
Model Fine-tuned MAE RMSE
BMNet [18] 8.05 9.7
BMNet+ [18] 5.76 7.83
CounTR [2] 5.75 7.45
BMNet [18] $\times$ 14.61 24.60
BMNet+ [18] $\times$ 10.44 13.77
LOCA [21] $\times$ 9.97 12.51
MGCAC (Ours) $\times$ 6.41 8.81
@@ 6.3 Ablation Studies
We perform ablation studies on MGCAC to examine the functionality of our proposed framework in the FSC-147 dataset and mosaic dataset. Based on the results, we have the following observations:
Mosaic Augmentation. In Table 5(a), comparing S1 and S2, the introduction of mosaic augmentation (S2) worsens the performance on FSC-147 dataset. In contrast, S2 gets significant improvement on the mosaic dataset (Table 5(b)). From Figure 7(c), we observe that mosaic augmentation allows S2 to focus on the correct region. Nevertheless, S2 still assigns density values on irrelevant objects and gives inaccurate counts.
Generalized Loss. With only applying GL (comparing S1 and S3), although S3 gets huge improvement on FSC-147 dataset (Table 5(a)), its performance becomes worse on mosaic dataset (Table 5(b)). As shown in Figure 7(d), although S3 captures the target objects, it also includes irrelevant objects. These observations indicate that while the use of GL enhances the localization ability, it inadvertently leads to model overfitting in single-class settings, hindering its effectiveness in differentiating classes.
@@
Multi-scale Feature with U-Net-like Fusion. Comparing S4 and S5 quantitatively, the incorporation of multi-scale feature aggregation (i.e. the multi-scale features from the feature extractor and the matcher are fused via a U-Net-like manner in the density head) allows MGCAC to be competitive on the FSC-147 dataset while still demonstrating its ability to distinguish reference classes on the mosaic dataset. The predicted count closely aligns with the ground truth in Figure 7(f).
6.4 Impact of Number of References Images (i.e. $n$)
We investigate our model performance in relation to the number of reference images, where $n\leq 3$. In Table 6, the results indicate that using our training protocol, our robust model architecture effectively captures reference representative features and improves performance as the number of reference images increases.
Table 5: The ablation results of the proposed MGCAC on FSC-147 and mosaic dataset, where ‘MA’ denotes using mosaic augmentation, ‘GL’ denotes replacing the MSE loss with the generalized loss, ‘MS’ means that the multi-scale features from the feature extractor and the matcher are fused via a U-Net-like manner in the density head.
@@
(b) Ablation results on mosaic dataset.
@@
Table 6: Evaluation results of MGCAC using different numbers (i.e. $n$) of reference images on FSC-147.
$n$ VAL TEST
MAE RMSE NAE SRE MAE RMSE NAE SRE
1 17.78 68.9 0.19 3.05 14.55 111.04 0.16 2.79
2 12.80 58.83 0.13 2.34 11.03 107.09 0.15 4.24
3 11.00 51.42 0.12 2.05 10.46 96.60 0.16 6.17

### v2

@@ Abstract
Class agnostic counting (CAC) is a vision task that can be used to count the total occurrence number of any given reference objects in the query image. The task is usually formulated as a density map estimation problem through similarity computation among a few image samples of the reference object and the query image. In this paper, we point out a severe issue of the existing CAC framework: Given a multi-class setting, models don’t consider reference images and instead blindly match all dominant objects in the query image. Moreover, the current evaluation metrics and dataset cannot be used to faithfully assess the model’s generalization performance and robustness. To this end, we discover that the combination of mosaic augmentation with generalized loss is essential for addressing the aforementioned issue of CAC models to count objects of majority (i.e. dominant objects) regardless of the references. Furthermore, we introduce a new evaluation protocol and metrics for resolving the problem behind the existing CAC evaluation scheme and better benchmarking CAC models in a more fair manner. Besides, extensive evaluation results demonstrate that our proposed recipe can consistently improve the performance of different CAC models. The code is available at https://github.com/littlepenguin89106/MGCAC.
@@ 1 Introduction
Through in-depth analysis, we pinpoint the root causes of RO problem: the lack of variety of classes in each query image during the training stage incapacitates the model’s need to differentiate among object classes. Additionally, using pixel-wise losses (e.g., MSE) fails to accurately localize target objects and still contributes to the model’s inability to differentiate objects effectively. To address these challenges, we propose the Mosaic-based Generalized-Loss (MG) recipe for CAC, which consistently allows different CAC models to effectively capture variations in reference appearances and precisely localize target objects as illustrated in Figure 1(b) and demonstrated by our extensive experimental results. Finally, we also provide a strong CAC baseline by exploiting more advanced cross-attention modules from [5] as the backbone with our proposed MG loss recipe for training, denoted as MGCAC. The proposed MGCAC achieves the state-of-the-art results for CAC upon the original FSC-147 [17] and our proposed mosaic evaluation dataset.
@@ 3.3 Benchmarking
Furthermore, the current CAC evaluation protocol evaluates model counting performance by aggregating counting errors across all images. Given groundtruth count $c_{l}$ and predicted count $\tilde{c_{l}}$ for the $l$th image in the dataset with $L$ images, $\text{MAE}=\frac{1}{L}\sum_{i=1}^{L}|\tilde{c_{l}}-c_{l}|$ and $\text{RMSE}=\sqrt{\frac{1}{L}\sum_{i=1}^{L}(\tilde{c_{l}}-c_{l})^{2}}$ are usually used. The challenge arises from the long-tailed distribution of per-image object counts, introducing bias towards large-count query images. Figure 4 illustrates that only two images featuring large counts in FSC-147 test dataset. By excluding the first two highest object-count query images from RMSE calculations on the FSC-147 test dataset (Table 1(b)), we show significant impacts on the evaluation results (details in Appendix). Recently SOTA models (i.e., LOCA and CounTR) get worse RMSE than BMNet+ when removing large-count queries. Those phenomena reinforce the need for an evaluation strategy that accounts for the distributional characteristics of object counts, especially in datasets with varying count sizes.
@@ 4.1 Mosaic Evaluation Dataset
Currently, there is a lack of real-world multi-class counting datasets for CAC evaluation. We leverage the FSC-147 dataset as a foundation to create a multi-class setting. As shown in Figure 5, for a synthesized mosaic collage, we randomly select 4 query images of distinct classes from the validation and test sets of FSC-147. These images are then randomly cropped to the size of 384 × 384 and put together as size 768 × 768. Each class in the collage will individually serve as the reference class. Given the associated reference images, the model has to accurately predict the number of target objects for the collage (i.e. predicted count should equate to the ground truth count of the query image containing the target class). In total, we generate 1,000 queries (i.e., 4,000 data pairs), facilitating a comprehensive evaluation of the object recognition capabilities within the context of mosaic collages. We name this dataset as FSC-Mosaic, which stands for Few-Shot Counting Dataset with Mosaic Evaluation.
@@ 5.2 Generalized Loss
Currently, most CAC approaches based on density map estimation employ the pixel-wise $L_{2}$ loss (or MSE loss) to train the models where the ground truth density map is generated by converting the point annotations of each training image through convolving with a Gaussian kernel. However, $L_{2}$ loss only offers pixel-wise supervision. It penalizes the same no matter whether the predictions are near or far from the ground-truth dot annotation. This makes $L_{2}$ loss inappropriate for accurate localization. Recently, Wan et al. [22] propose a Generalized Loss (GL) that uses unbalanced optimal transport (OT) to measure the transport cost between the predicted density map and the ground truth point annotations directly. GL penalizes more when the predictions are farther from the ground truth. This suffices the need for precise localization for single and multi-class settings. In addition, it also has shown that the $L_{2}$ loss is a special case of GL and usually results in suboptimal solutions. Thus, we mainly exploit GL to train the proposed models shown as follows:
@@ 6.1 Implementation Details
Datasets. We utilize three datasets to evaluate our model performance: (1) FSC-147 dataset [17], a standard class-agnostic object counting dataset composed of 147 categories and 6,135 images across different scales, in which each image is annotated with three reference objects. Given that the dataset contains decent annotations, we use the FSC-147 dataset to train and evaluate our models. (2) We evaluate FSC-147-trained models on our proposed FSC-Mosaic dataset to validate their generalizability in multi-class scenarios. (3) Moreover, we evaluate FSC-147-trained models on CARPK [7], a dataset that contains 1,448 photos of cars in parking lots from a bird-view perspective, to ensure the generalizations of our models on unseen datasets (see Appendix).
@@
Model Comparison. We utilize MixFormer [5] architecture design with a multi-scale density head to better localize features of different levels as our baseline, denoted as MGCAC as mentioned in Section 1. (Please refer to Appendix for more detailed descriptions.) It is worth noting that our training recipe can be applied to other CAC model architectures as well. To demonstrate the effectiveness and general applicability of our proposed training recipe, we compare BMNet+, CounTR, and LOCA with their original training approach against our proposed recipe on the FSC-Mosaic dataset.
@@ 6.2 Evaluation Results with Other SOTAs
Mosaic Evaluation. Seen from Table 3, models under the CAC typical training process, such as BMNet+ and LOCA, exhibit unsatisfactory performance under mosaic evaluation. The introduction of mosaic augmentation, as demonstrated by CounTR, yields a better performance. After employing our training recipe, BMNet+, CounTR, and LOCA exhibit substantial error decreases in all metrics. Specifically, BMNet+ yields an improvement of 47.5% in RMSE and over 75% across the remaining three metrics. These improvements demonstrate the importance and efficacy of applying the combination of mosaic augmentation and generalized loss on CAC models in discriminating reference classes and accurately matching targets. Notably, MGCAC achieves SOTA on the FSC-Mosaic dataset with 37% improvement on all metrics against all models, showing our superiority in counting target objects.
@@
Table 3: Evaluation results on the FSC-Mosaic dataset of the CAC models trained with the proposed training recipe (Red) against those using their original training recipe where we use their publicly available pre-trained models (Green).
@@ 6.3 Ablation Studies
We perform ablation studies on MGCAC to examine the functionality of our proposed framework in the FSC-147 dataset and FSC-Mosaic dataset. Based on the results, we have the following observations:
Mosaic Augmentation. In Table 4(a), comparing S1 and S2, the introduction of mosaic augmentation (S2) worsens the performance on FSC-147 dataset. In contrast, S2 gets significant improvement on the FSC-Mosaic dataset (Table 4(b)). From Figure 6(c), we observe that mosaic augmentation allows S2 to focus on the correct region. Nevertheless, S2 still assigns density values on irrelevant objects and gives inaccurate counts.
Generalized Loss. With only applying GL (comparing S1 and S3), although S3 gets huge improvement on FSC-147 dataset (Table 4(a)), its performance becomes worse on FSC-Mosaic dataset (Table 4(b)). As shown in Figure 6(d), although S3 captures the target objects, it also includes irrelevant objects. These observations indicate that while the use of GL enhances the localization ability, it inadvertently leads to model overfitting in single-class settings, hindering its effectiveness in differentiating classes.
@@
Multi-scale Feature with U-Net-like Fusion. Comparing S4 and S5 quantitatively, the incorporation of multi-scale feature aggregation (i.e. the multi-scale features from the feature extractor and the matcher are fused via a U-Net-like density head) allows MGCAC to be competitive on the FSC-147 dataset while still demonstrating its ability to distinguish reference classes on the FSC-Mosaic dataset. The predicted count closely aligns with the ground truth in Figure 6(f).
Table 4: The ablation results of the proposed MGCAC on FSC-147 and FSC-Mosaic dataset, where ‘MA’ denotes using mosaic augmentation, ‘GL’ denotes replacing the MSE loss with the generalized loss, ‘MS’ means that the multi-scale features from the feature extractor and the matcher are fused via a U-Net-like density head.
@@
(b) Ablation results on FSC-Mosaic dataset.
@@ 7 Conclusion
8 Acknowledgements
This research work is supported in part by National Science and Technology Council (NSTC) under grants 112-$2$-E-001-001-MY2, 112-2634-F-002-006, 112-2221-E-A49-087-MY3, 111-2628-E-A49-018-MY4, and by Academia Sinica under grant AS-CDA-110-M09. We also thank to National Center for High-performance Computing (NCHC) of National Applied Research Laboratories (NARLabs) in Taiwan for providing computational and storage resources.